OpenStudy (anonymous):

can someone help me with Nature of roots and zeros? Find all the zeros of each polynomial function. f(x) = x4 – 5x2 – 36 1) Select one: a. ±3, ±2i b. ±3 only c. ±2i only d. none of the above

OpenStudy (anonymous):

To find the roots of a polynomial equation you should: Select one: a. only use Descartes Rule of Signs. b. only use the Quadratic Formula. c. only factor it. d. use any method that gives the roots of the function.

OpenStudy (barrycarter):

Hint (for original problem): substitute y = x^2

OpenStudy (anonymous):

huh?

OpenStudy (campbell_st):

what has been suggested is that this equation is reducible to a quadratic equation... this can be done by using an appropriate substitution... such as letting y = x^2 so the equation becomes \[y^2 - 5y - 36 \] now you can use the general quadratic formula \[y = \frac{5 \pm \sqrt{25 - 144}}{2}\] then when you get the values for y, replace it with x^2 and then solve for x